Calculate what happens to nominal GDP if velocity remains constant at 4 and the money supply increases 50%. Assume that GDP is now at $2 trillion
1
Expert's answer
2020-12-10T07:52:52-0500

Constant velocity = 4

Nominal GDP1 = $2 trillion

"Velocity = \frac{Nominal \;GDP}{Money \;supply} \\

4 = \frac{2\; trillion}{Money \;supply_1} \\

Money \;supply_1 = 0.5 \; trillion \\

Money \;supply_2 = Money \;supply_1 + Money \;supply_1 \times 0.5 \\

= 0.5 + 0.25 = 0.75 \; trillion \\

4 = \frac{Nominal \;GDP_2}{0.75 \; trillion} \\

Nominal \;GDP_2 = 3 \; trillion"

Nominal GDP will increase from $2 trillion to $3 trillion.